# One Prompt To Rule Them All: LLMs for Opinion Summary Evaluation

### v1

@@ Abstract
Evaluation of opinion summaries using conventional reference-based metrics rarely provides a holistic evaluation and has been shown to have a relatively low correlation with human judgments. Recent studies suggest using Large Language Models (LLMs) as reference-free metrics for NLG evaluation, however, they remain unexplored for opinion summary evaluation. Moreover, limited opinion summary evaluation datasets inhibit progress. To address this, we release the SummEval-Op dataset covering $7$ dimensions related to the evaluation of opinion summaries: fluency, coherence, relevance, faithfulness, aspect coverage, sentiment consistency, and specificity. We investigate Op-I-Prompt, a dimension-independent prompt, and Op-Prompts, a dimension-dependent set of prompts for opinion summary evaluation. Experiments indicate that Op-I-Prompt emerges as a good alternative for evaluating opinion summaries achieving an average Spearman correlation of $\mathbf{0.70}$ with humans, outperforming all previous approaches. To the best of our knowledge, we are the first to investigate LLMs as evaluators on both closed-source and open-source models in the opinion summarization domain.
One Prompt To Rule Them All: LLMs for Opinion Summary Evaluation
Tejpalsingh $*{}^{*}$${}^{\clubsuit}$, Swaroop $*{}^{*}$${}^{\clubsuit}$, Sri $*{}^{*}$${}^{\clubsuit}$, Rupasai $*{}^{*}$${}^{\clubsuit}$, Swaprava ${}^{\clubsuit}$, Pushpak ${}^{\clubsuit}$, Suman ${}^{\spadesuit}$, Amey ${}^{\spadesuit}$, Sudhanshu Shekhar ${}^{\spadesuit}$, Muthusamy ${}^{\spadesuit}$, Nikesh ${}^{\spadesuit}$$^{\clubsuit}$omputer Science and Engineering, IIT Bombay, India, ${}^{\spadesuit}$lipkart, India {tejpalsingh, swaroopnath, sriraghava, rupasai, swaprava, pb}@cse.iitb.ac.in
**footnotetext: Equal contribution
@@ 1 Introduction
    SummEval-Op***https://github.com/tjsiledar/SummEval-OP, an opinion summarization benchmark dataset, consisting of a total of $2,912$ summary annotations, assessing $13$ opinion summaries for $32$ products from the Amazon test set. The evaluation covers $\mathbf{7}$ dimensions- fluency, coherence, relevance, faithfulness, aspect coverage, sentiment consistency, and specificity related to the evaluation of opinion summaries (Section 4).
@@
    Benchmarking of recent LLMs (closed and open-source) on the aforementioned $7$ dimensions for the task of opinion summarization, which to the best of our knowledge is first of its kind (Section 6).
@@
    Detailed analysis, comparing an open-source LLM against a closed-source LLM acting as evaluators for automatic evaluation of opinion summaries on $7$ dimensions. Analysis indicates that Op-I-Prompt emerges as a good alternative for evaluating opinion summaries showing a high correlation with humans when compared with alternatives (Section 6).
@@ 2 Related Work
LLM-based EvaluatorsFu et al. (2023) introduced GPTScore that operates on the premise that a generative pre-training model (e.g. GPT-3) is likely to assign a higher probability to the generation of high-quality text in line with provided instructions and context. Chiang and Lee (2023a) were the first to explore LLMs for evaluation. Chiang and Lee (2023b) provide concrete guidelines that improve ChatGPT’s correlation with humans. Wang et al. (2023) conducted an initial survey exploring the utilization of ChatGPT as an NLG evaluator. Kocmi and Federmann (2023) used GPT models for evaluating machine learning tasks. Liu et al. (2023) introduced G-Eval, a framework for evaluation of NLG outputs using Chain of Thought (CoT) (Wei et al., 2023) and assigning weights to a predetermined set of integer scores based on their generation probabilities from GPT-3/4. Chen et al. (2023) were the first to investigate approaches to reference-free NLG evaluation using LLMs, finding that an explicit score generated by ChatGPT is the most effective and stable approach. Zheng et al. (2023) show that strong LLMs such as GPT-4 achieve a similar level of agreement to that of humans and hence can be used to approximate human preferences. Our work investigates two prompt strategies and tests the applicability of different prompt approaches on closed-source and open-source LLMs for opinion summary evaluation for $7$ dimensions.
Opinion Summary Evaluation Benchmark Shen and Wan (2023) created the OpinSummEval dataset, utilizing the Yelp test set (Chu and Liu, 2019), annotating for $4$ dimensions relevant to opinion summary evaluation. Our work enhances this effort by introducing SummEval-OP, which focuses on the e-commerce domain, constructed using the Amazon test set Bražinskas et al. (2020). Additionally, we collect annotations for $7$ dimensions on the recent LLM summaries, subsequently establishing benchmarks for comparison.
@@ 3.1 Prompt Approaches
Task Description: Defines the task that the LLM will be performing. In our case, the task is to evaluate a summary corresponding to a set of reviews on a given metric/dimension.
Evaluation Criteria: Defines the criteria that will be used to perform the task. In our case, the task being opinion summary evaluation, the criteria is to assign a score ($1-5$) for a certain metric/dimension depending on the extent to which the summary adheres to it.
Evaluation Steps: This comprises the steps that the LLM must take to correctly perform the described task. In our case, it contains the steps that the LLM should follow to evaluate a certain metric/dimension.
@@ 4.1 Dataset
We utilized the Amazon test set (He and McAuley, 2016; Bražinskas et al., 2020), comprising of reviews from $4$ domains: electronics, home & kitchen, personal care, and clothing, shoes & jewelry. The test set contained a total of $32$ products, each with $3$ human-annotated reference summaries and $8$ reviews per product. For our use, we needed one human reference summary per product which we obtained by randomly selecting one of the summaries out of the $3$ for each product. We do not directly consider only one of the human summaries as this would bias the summaries to a single person.
@@ 4.3 Annotation Details
For creating the SummEval-Op dataset, annotations were collected for a total of $13$ summaries per product across $7$ dimensions for $32$ products from the Amazon test set. The $13$ summaries comprised of $1$ human-annotated reference summary (as mentioned in Section 4.1) and $12$ different model-generated summaries (listed in Section 5.1). To ensure high quality of annotations, each summary was annotated by $3$ raters for $7$ dimensions, amounting to $2,912$ summary-level ratings. Raters were asked to rate the summaries on a Likert scale from $1$ to $5$ (higher is better) along the $7$ dimensions. Each summary for each dimension was rated by $3$ raters. The overall quantity of annotations is: $3$ (# of raters) x $32$ (# of instances) x $13$ (# of summaries) x $7$ (# of dimensions) = $8,736$ ratings.
We chose to hire $3$ Masters’ students with experience in opinion summarization as opposed to crowd workers for the following reasons: (a) Gillick and Liu (2010) demonstrated that summary evaluations from non-experts can significantly diverge from expert annotations and may display inferior inter-annotator agreement, and (b) Fabbri et al. (2021) emphasized the significance of employing expert annotators to mitigate quality concerns in human ratings. Similar to Fabbri et al. (2021), we conducted the process in two rounds, to ensure high-quality ratings. In Round II, ratings where the scores of any rater differed from any other rater by $2$ or more points were re-evaluated. The re-evaluation was done through a discussion between the annotators such that ratings from all $3$ differ by at most $1$. We asked the raters to be critical and discuss the ratings during re-evaluation. Check Appendix B
@@ 4.4 Annotation Analysis
We evaluated the inter-rater agreement for the $3$ raters using Krippendorff’s alpha coefficient ($\alpha$) (Krippendorff, 2011). For Round-I, we found the coefficient to be $0.50$ indicating moderate aggrement ($0.41\leq\alpha\leq 0.60$). For Round-II, the coefficient increased to $0.80$, indicating substantial agreement ($0.61\leq\alpha\leq 0.80$). We report the dimension-wise agreement scores for both rounds in Table 1. We observe that for both Round-I and Round-II, faithfulness and aspect coverage score higher than others. This is mostly because faithfulness and aspect coverage could be identified by cross-examining with the reviews. After Round-II, coherence and specificity are the most disagreed upon between raters. This could be attributed to their subjective nature (Kryściński et al., 2018).
@@ 5 Experiments
UniEval 0.36 0.28 0.52 0.42 0.33 0.25 0.17 0.14 - - - - - -
G-Eval-3.5 0.63 0.55 0.59 0.49 0.68 0.56 0.70 0.58 0.79 0.67 0.73 0.61 0.75 0.63
OP-I-GPT-3.5 0.60 0.51 0.61 0.51 0.69 0.56 0.71 0.59 0.80 0.68 0.73 0.61 0.74 0.61
@@
OP-I-Mistral 0.54 0.45 0.58 0.47 0.59 0.47 $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
Table 2: Spearman ($\rho$) and Kendall Tau ($\tau$) correlations at summary-level on $7$ dimensions for the SummEval-OP dataset. For closed-source, Op-I-Prompt performs comparably to G-Eval, whereas for open-source it outperforms alternatives. *** represents significant performance (p-value $<0.05$) to G-Eval-Mistral computed using Mann-Whitney U Test. Humans- averaged correlation of each annotator with the overall averaged ratings.
5.1 Summarization Models
Pre-LLMs: For the Pre-LLMs, we obtain the publicly available summaries for the Amazon test set of these models. These models were trained in a self-supervised manner using only reviews data. (1) PlanSum (Amplayo and Lapata, 2020) uses content plans to create relevant review-summary pairs. The content plans take the form of aspect and sentiment distributions which are used along with input reviews for generating summaries. (2) MultimodalSum (Im et al., 2021) uses non-text data such as image and metadata along with reviews to generate opinion summaries. It uses a separate encoder for each modality and uses synthetic datasets to train the model in an end-to-end fashion. (3) Siledar et al. (2023) uses lexical and semantic similarities to create a highly relevant synthetic dataset of review-summary pairs. This is then used to fine-tune any pre-trained language model for generating opinion summaries. (hereby referred to as LS-Sum-G).
[…]
5.2 Baselines
5.3 Implementation Details
For evaluation, we used Mistral-$7$B (mistralai/Mistral-7B-Instruct-v0.2) as our evaluator model. We chose Mistral-$7$B for these reasons: (a) it ranked best amongst the open-source models on the lmsys/chatbot-arena-leaderboard, (b) we found its instruction following-ness to be better than alternatives, and (c) its $7$B size ensures easy replication. We set the hyperparameters to n=100, temperature=0.7 to sample multiple generations. Example prompts are in Appendix D.
@@ 5.3 Implementation Details
ChatGPT-$3.5$ 4.89 4.58 4.25 4.71 4.22 4.16 3.96
GPT-$4$ 5.00 4.91 3.52 4.96 4.93 4.83 4.57
@@
LLaMA2222-$13$B 4.87 4.49 4.25 4.62 4.02 4.00 3.94
Mistral-$7$B 4.86 4.60 4.33 4.66 4.56 4.35 4.25
@@
Solar-$10.7$B 4.89 4.73 4.20 4.72 4.50 4.56 4.35
Zephyr-$7$B 4.89 4.36 4.08 4.54 4.18 3.95 3.83
Table 3: Model-wise averaged annotator ratings of opinion summaries along $7$ dimensions for the Amazon test set. Best scores are in bold, second-best are underlined.

### v2

# Sudhanshu Shekhar Singh, Muthusamy Chelliah, Nikesh Garera Computer Science and Engineering, IIT Bombay, India, Flipkart, India

@@ Abstract
Evaluation of opinion summaries using conventional reference-based metrics often fails to provide a comprehensive assessment and exhibits limited correlation with human judgments. While Large Language Models (LLMs) have shown promise as reference-free metrics for NLG evaluation, their potential remains unexplored for opinion summary evaluation. Furthermore, the absence of sufficient opinion summary evaluation datasets hinders progress in this area. In response, we introduce the SummEval-Op dataset, encompassing $7$ dimensions crucial to the evaluation of opinion summaries: fluency, coherence, relevance, faithfulness, aspect coverage, sentiment consistency, and specificity. We propose Op-I-Prompt, a dimension-independent prompt, along with Op-Prompts, a dimension-dependent set of prompts for opinion summary evaluation. Our experiments demonstrate that Op-I-Prompt emerges as a good alternative for evaluating opinion summaries, achieving an average Spearman correlation of $\mathbf{0.70}$ with human judgments, surpassing prior methodologies. Remarkably, we are the first to explore the efficacy of LLMs as evaluators, both on closed-source and open-source models, in the opinion summary evaluation domain.
[Uncaptioned image]ne Prompt To Rule Them All: LLMs for Opinion Summary Evaluation
Tejpalsingh Siledar, Swaroop Nath, Sri Raghava, Rupasai Rangaraju, Swaprava Nath, Pushpak Bhattacharyya, Suman Banerjee, Amey Patil, Sudhanshu Shekhar Singh, Muthusamy Chelliah, Nikesh Garera Computer Science and Engineering, IIT Bombay, India, Flipkart, India {tejpalsingh, swaroopnath, sriraghava, rupasai, swaprava, pb}@cse.iitb.ac.in
Equal contribution
@@ 1 Introduction
    SummEval-Op [^1] https://github.com/tjsiledar/SummEval-OP, an opinion summary evaluation benchmark dataset, consisting of a total of $2,912$ summary annotations, assessing $13$ opinion summaries for $32$ products from the Amazon test set. The evaluation covers $\mathbf{7}$ dimensions- fluency, coherence, relevance, faithfulness, aspect coverage, sentiment consistency, and specificity related to the evaluation of opinion summaries (Section 4).
@@
    Benchmarking of $\mathbf{9}$ recent LLMs (closed and open-source) on the aforementioned $7$ dimensions for the task of opinion summarization, which to the best of our knowledge is first of its kind (Table 4, Section 6).
@@
    Detailed analysis, comparing an open-source LLM against a closed-source LLM acting as evaluators for automatic evaluation of opinion summaries on $7$ dimensions. Analysis indicates that Op-I-Prompt emerges as a good alternative for evaluating opinion summaries showing a high correlation (spearman correlation of $\mathbf{0.70}$ on average) with humans when compared with alternatives (Section 6).
@@ 2 Related Work
LLM-based Evaluators.Fu et al. (2023) introduced GPTScore that operates on the premise that a generative pre-training model (e.g. GPT-3) is likely to assign a higher probability to the generation of high-quality text in line with provided instructions and context. Chiang and Lee (2023a) were the first to explore LLMs for evaluation. Chiang and Lee (2023b) provide concrete guidelines that improve ChatGPT’s correlation with humans. Wang et al. (2023) conducted an initial survey exploring the utilization of ChatGPT as an NLG evaluator. Kocmi and Federmann (2023) used GPT models for evaluating machine learning tasks. Liu et al. (2023) introduced G-Eval, a framework for evaluation of NLG outputs using Chain of Thought (CoT) (Wei et al., 2023) and assigning weights to a predetermined set of integer scores based on their generation probabilities from GPT-3/4. Chen et al. (2023) were the first to investigate approaches to reference-free NLG evaluation using LLMs, finding that an explicit score generated by ChatGPT is the most effective and stable approach. Zheng et al. (2023) show that strong LLMs such as GPT-4 achieve a similar level of agreement to that of humans and hence can be used to approximate human preferences. Our work investigates two prompt strategies and tests the applicability of different prompt approaches on closed-source and open-source LLMs for opinion summary evaluation for $7$ dimensions.
Opinion Summary Evaluation Benchmark. Shen and Wan (2023) created the OpinSummEval dataset, utilizing the Yelp test set (Chu and Liu, 2019), annotating for $4$ dimensions relevant to opinion summary evaluation. Our work enhances this effort by introducing SummEval-Op, which focuses on the e-commerce domain, constructed using the Amazon test set Bražinskas et al. (2020). Additionally, we collect annotations for $7$ dimensions on the recent LLM summaries, subsequently establishing benchmarks for comparison.
Opinion Summarization  Opinion summarization aims to summarize opinions into concise summaries (Wang and Ling, 2016; Siledar et al., 2023b). Recent approaches use unsupervised methods (Chu and Liu, 2019; Bražinskas et al., 2020), or creating synthetic datasets with pseudo-summaries (Bražinskas et al., 2020; Amplayo and Lapata, 2020; Amplayo et al., 2021; Elsahar et al., 2021; Im et al., 2021; Siledar et al., 2023a, 2024). Notably, Amplayo et al. (2021) and Im et al. (2021) advanced these methods by generating content plans and using multimodal inputs, respectively. Siledar et al. (2023a) used lexical and semantic similarities to generate synthetic data. In this work, we benchmark the recent LLMs for the task of opinion summarization.
@@ 3.1 Prompt Approaches
(1) Task Description: Defines the task that the LLM will be performing. In our case, the task is to evaluate a summary corresponding to a set of reviews on a given metric/dimension.
(2) Evaluation Criteria: Defines the criteria that will be used to perform the task. In our case, the task being opinion summary evaluation, the criteria is to assign a score ($1-5$) for a certain metric/dimension depending on the extent to which the summary adheres to it.
(3) Evaluation Steps: This comprises the steps that the LLM must take to correctly perform the described task. In our case, it contains the steps that the LLM should follow to evaluate a certain metric/dimension.
@@
3.2 Prompt Design Consideration
The design of our prompts was based on the intuition that LLMs would produce improved responses when prompted to justify their evaluations. Our approach ensures that the response reiterates the evaluation metric, highlights both strengths and shortcomings and concludes with an evaluation score based on the criteria outlined in the prompt.
@@ 4.1 Dataset
We utilized the Amazon test set (He and McAuley, 2016; Bražinskas et al., 2020), comprising of reviews from $4$ domains: electronics, home & kitchen, personal care, and clothing, shoes & jewelry. The test set contained a total of $32$ products, each with $3$ human-annotated abstractive summaries and $8$ reviews per product. The reviews and summaries are all in English. For our use, we needed one human reference summary per product which we obtained by randomly selecting one of the summaries out of the $3$ for each product. We do not directly consider only one of the human summaries as this would bias the summaries to a single person.
@@ 4.3 Annotation Details
For creating the SummEval-Op dataset, annotations were collected for a total of $13$ abstractive summaries per product across $7$ dimensions for $32$ products from the Amazon test set. The $13$ summaries comprised of $1$ human-annotated reference summary (mentioned in Section 4.1) and $12$ different model-generated summaries (listed in Section 5.1). To ensure high quality of annotations, each summary was annotated by $3$ raters for $7$ dimensions, amounting to $2,912$ summary-level ratings. Raters were asked to rate the summaries on a scale from $1$ to $5$ (higher is better) along the $7$ dimensions. Each summary for each dimension was rated by $3$ raters. The overall quantity of annotations is: $3$ (# of raters) x $32$ (# of instances) x $13$ (# of summaries) x $7$ (# of dimensions) = $8,736$ ratings.
We hired $3$ Masters’ students with experience in opinion summarization as opposed to crowd workers for the following reasons: (a) Gillick and Liu (2010) demonstrated that summary evaluations from non-experts can significantly diverge from expert annotations and may display inferior inter-annotator agreement, and (b) Fabbri et al. (2021) emphasized the significance of employing expert annotators to mitigate quality concerns in human ratings. Similar to Fabbri et al. (2021), we conducted the process in two rounds, to ensure high-quality ratings. In Round II, ratings where the scores of any rater differed from any other rater by $2$ or more points were re-evaluated. The re-evaluation was done through a discussion between the annotators such that ratings from all $3$ differ by at most $1$. We asked the raters to be critical and discuss the ratings during re-evaluation.
We hired raters who have written papers on opinion summarization (1) or are working in the opinion summarization domain (2). These were male Masters’ students aged 21-30. They were provided with detailed guidelines for evaluating summaries on the $7$ dimensions. All $3$ raters received stipends suitable for the tasks. Models associated with summaries were not revealed to the raters to remove any bias. Check Appendix B.
@@ 4.4 Annotation Analysis
We evaluated the inter-rater agreement for the $3$ raters using Krippendorff’s alpha coefficient ($\alpha$) (Krippendorff, 2011). For Round-I, we found the coefficient to be $0.50$ indicating moderate aggrement ($0.41\leq\alpha\leq 0.60$). For Round-II, the coefficient increased to $0.80$, indicating substantial agreement ($0.61\leq\alpha\leq 0.80$). We report the dimension-wise agreement scores for both rounds in Table 1. We observe that for both Round-I and Round-II, faithfulness and aspect coverage score higher than others. This is mostly because faithfulness and aspect coverage could be identified by cross-examining with the reviews. After Round-II, coherence and specificity are the most disagreed upon between raters.
@@
UniEval $0.36$ $0.28$ $0.52$ $0.42$ $0.33$ $0.25$ $0.17$ $0.14$- -- -- -- -- --
G-Eval-3.5 $\mathbf{0.63}$ $\mathbf{0.55}$ $0.59$ $\underline{0.49}$ $\underline{0.68}$ $\mathbf{0.56}$ $\underline{0.70}$ $\underline{0.58}$ $0.79$ $0.67$ $\mathbf{0.73}$ $\mathbf{0.61}$ $\mathbf{0.75}$ $\mathbf{0.63}$
OP-I-GPT-3.5 $\underline{0.60}$ $\underline{0.51}$ $\mathbf{0.61}$ $\mathbf{0.51}$ $\mathbf{0.69}$ $\mathbf{0.56}$ $\mathbf{0.71}$ $\mathbf{0.59}$ $\underline{0.80}$ $\underline{0.68}$ $\mathbf{0.73}$ $\mathbf{0.61}$ $\underline{0.74}$ $\underline{0.61}$
@@
OP-I-Mistral $0.54$ $0.45$ $0.58$ $0.47$ $0.59$ $0.47$ $0.63^{*}$ $0.51^{*}$ $\mathbf{0.82}^{*}$ $\mathbf{0.70}^{*}$ $\mathbf{0.73}^{*}$ $\mathbf{0.61}^{*}$ $0.71^{*}$ $0.58^{*}$
Table 2: Spearman ($\rho$) and Kendall Tau ($\tau$) correlations at summary-level on $7$ dimensions for the SummEval-Op dataset (Round-I). For closed-source, Op-I-Prompt performs comparably to G-Eval, whereas for open-source it outperforms alternatives. * represents significant performance (p-value $<0.05$) to G-Eval-Mistral computed using Mann-Whitney U Test. Humans- averaged correlation of each annotator with the overall averaged ratings.
G-Eval Response
<score>5</score> Explanation: The summary covers all the major aspects discussed in the reviews, including the comfort and break-in period of the boots. It also mentions that the boots take a little while to break in, which is a common theme among the reviews. Therefore, the summary receives a high rating for its aspect coverage.
Op-I-Prompt Response
The metric for evaluation is Aspect Coverage. The summary should cover all the aspects that are majorly being discussed in the reviews. The reviews mention the following aspects: comfort and break-in period. The summary states "I love Frye boots. They are the most comfortable boots I have ever worn." This covers the comfort aspect. The summary also states "They do take a little while to break in, but they are worth the wait." This covers the break-in period aspect. However, the summary does not mention anything about the durability, fit and appearance aspects. Since these aspects were majorly being discussed in the reviews, the summary does not follow the metric completely. Score- <score>2</score> The summary only follows the metric to a limited extent while generating the summary from the reviews.
Table 3: Comparative Analysis. Sample responses generated by our Op-I-Prompt compared to G-Eval for evaluating a sample summary for the dimension aspect coverage. On average, the scores assigned for this summary by annotators, G-Eval, and Op-I-Prompt are $2.33$, $4.09$, and $2.72$ respectively.
@@
ChatGPT-$3.5$ $\underline{4.89}$ $4.58$ $\mathbf{4.25}$ $4.71$ $4.22$ $4.16$ $3.96$
GPT-$4$ $\mathbf{5.00}$ $\mathbf{4.91}$ $3.52$ $\mathbf{4.96}$ $\mathbf{4.93}$ $\mathbf{4.83}$ $\mathbf{4.57}$
@@
LLaMA2222-$13$B $4.87$ $4.49$ $\mathbf{4.25}$ $4.62$ $4.02$ $4.00$ $3.94$
Mistral-$7$B $4.86$ $4.60$ $\underline{4.33}$ $4.66$ $\underline{4.56}$ $4.35$ $4.25$
@@
Solar-$10.7$B $\underline{4.89}$ $\underline{4.73}$ $4.20$ $\underline{4.72}$ $4.50$ $\underline{4.56}$ $\underline{4.35}$
Zephyr-$7$B $\underline{4.89}$ $4.36$ $4.08$ $4.54$ $4.18$ $3.95$ $3.83$
Table 4: Model-wise averaged annotator ratings of opinion summaries along $7$ dimensions (Round-II). Best scores are in bold, second-best are underlined.
5 Experiments
5.1 Summarization Models
Pre-LLMs: For the Pre-LLMs, we obtain the publicly available summaries for the Amazon test set of these models. These models were trained in a self-supervised manner using only reviews data. (1) PlanSum (Amplayo et al., 2021) uses content plans to create relevant review-summary pairs. The content plans take the form of aspect and sentiment distributions which are used along with input reviews for generating summaries. (2) MultimodalSum (Im et al., 2021) uses non-text data such as image and metadata along with reviews to generate opinion summaries. It uses a separate encoder for each modality and uses synthetic datasets to train the model in an end-to-end fashion. (3) Siledar et al. (2023a) uses lexical and semantic similarities to create a highly relevant synthetic dataset of review-summary pairs. This is then used to fine-tune any pre-trained language model for generating opinion summaries. (hereby referred to as LS-Sum-G).
[…]
5.2 Baselines
5.3 Implementation Details
For evaluation, we used Mistral-$7$B (mistralai/Mistral-7B-Instruct-v0.2) as our evaluator model. We chose Mistral-$7$B for these reasons: (a) it ranked best amongst the open-source models on the lmsys/chatbot-arena-leaderboard, (b) we found its instruction following-ness to be better than alternatives, and (c) its $7$B size ensures easy replication. We set the hyperparameters to n=100, temperature=0.7 to sample multiple generations. Example prompts are in Appendix C.